OpenStudy (anonymous):

n - 10 = 5/6 n - 7 - 1/3 n

OpenStudy (anonymous):

multiple both sides by 6 6n -60=5n -42 -2n can you solve it further?

OpenStudy (anonymous):

6N - 60 = 3N - 42

OpenStudy (anonymous):

what is n ?

OpenStudy (anonymous):

question not clear,where does the divisor end

OpenStudy (anonymous):

and... i need help again

OpenStudy (anonymous):

n - 10 = (5/6) n - 7 - (1/3 n), is this what you mean it can be written many ways, write it again with paranthesis

OpenStudy (anonymous):

@heedcom it is just like that as you write

OpenStudy (anonymous):

what type of help @Rubie

OpenStudy (anonymous):

how do i simplify next ?

OpenStudy (anonymous):

@uzumakhi how did you know that without rubi telling us.

OpenStudy (anonymous):

yes take terms with N on left side and terms without N to right side

OpenStudy (anonymous):

I will leave this to you uzumakhi, as the question can be written many ways, and without rubie stating the brackets, then I cant solve

OpenStudy (anonymous):

is it like this? \[n-10=\frac{ 5 }{ 6n }-7-\frac{ 1 }{ 3n}\]

OpenStudy (anonymous):

@remainder thats what I have been telling rubie, but she does not reply with the proper question, she justs wants the answer, lol

OpenStudy (anonymous):

@heedcom i don't what does rubie means

OpenStudy (anonymous):

lol i'm sorry, i was doing my other questions, and @REMAINDER nope n is not there.

OpenStudy (anonymous):

oh nvm :) its okai.